7. The payback period The payback method helps firms establish and identify a maximum acceptable payback perod that helps in their capital budgeting dedisions. Consider the case of Cold Goose Metal Works Inc.: Cold Goose Metal Works Inc. is a small firm, and several of its managers are worried about how soon the firm will be able to recover its initial investment from Project Betas expected future cash flows. To answer this question, Cold Gooses CFO has asked that you compute the projedts payback period using the following expected net cash flows and assuming that the cash flows are received evenly throughout each year Complete the following table and compute the projects conventional payback period. For full credit, complete the entire table. Year o Year 1 Year 2 Year 3 Expected cash flow Cumulative cash flow -5,500,000 $2,200,000 $4,675,000 925,000 conventional payback period: The conventional payback period ignores the time value of money, and this conoerns Cold Gooses CrO. He has now asked you to compute Betes discounted payback perioe, assuming the ompany has a 10% cost of capital. Complete the following table and perform any necessary calculations. Round the discounted cash how values to the nearest whole dollar, and the discounted payback period to the nearest two decimal places. For full credit,complete the entire table. year o Year 1 Year 2 Year 3 Cash fow Discounted cash How Cumulative discounted cash ow -5,500,000$2,200,000$4,675,000$1,925,000 oisc inted paybeck perioe Which version of a projects paybeck period sheuls the cro use when evaluating Project Seta, given its thecretical superiority7 O The discounted payback period 0 The regular payback period one thecretical disodvantage of both payback methods- compared to the net present value method-is that they fail to consider the value of the cash Hows beyone the point i tme equal to the paybeck period How much value does the discounted payback pertos method tail to recognize due te ths thecretical defidency
Answer #1
Year0 Year1 Year2 Year3
Expected Cash Flow $ (5,500,000.00) $ 2,200,000.00 $ 4,675,000.00 $ 1,925,000.00
Cumulative Cash flow $ (5,500,000.00) $ (3,300,000.00) $ 1,375,000.00 $ 3,300,000.00
Conventional Payback Period 1.71 (=1+3300000/4675000)
Year0 Year1 Year2 Year3
Expected Cash Flow $ (5,500,000.00) $ 2,200,000.00 $ 4,675,000.00 $ 1,925,000.00
Discounted Cash Flow= Expected Cash Flow/(1+r)^n $ (5,500,000.00) $ 2,000,000.00 $ 3,863,636.36 $ 1,446,280.99
Cumulative Cash flow $ (5,500,000.00) $ (3,500,000.00) $    363,636.36 $ 1,809,917.36
Conventional Payback Period 1.75 (=1+3500000/3863636.36)

Discounted payback period is better.

Value which discounted payback period doesnot recognize = 3,863,636.36 + 1,446,280.89 = 5,309,917.36

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well